OpenStudy (james1769):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (michele_laino):

please keep in mind that the radical: \[\sqrt[3]{{{x^4}}}\] can be rewritten as below: \[{x^{4/3}}\]

OpenStudy (michele_laino):

it is the definition of powers with rational exponent

OpenStudy (michele_laino):

now, if I apply the rule of power of power, we can write this: \[\huge{\left( {\sqrt[3]{{{x^4}}}} \right)^5} = {\left( {{x^{4/3}}} \right)^5} = {x^{5 \cdot \frac{4}{3}}} = {x^{\frac{{20}}{3}}}\]

OpenStudy (michele_laino):

\[\Large {\left( {\sqrt[3]{{{x^4}}}} \right)^5} = {\left( {{x^{4/3}}} \right)^5} = {x^{5 \cdot \frac{4}{3}}} = {x^{\frac{{20}}{3}}}\]

OpenStudy (james1769):

okay thnks

OpenStudy (michele_laino):

:)